### v1

@@ 5.2 Results on Public Benchmark Data
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM Im-LS-$\mathfrak{U}$-QTSVM
pima $77.74\pm 2.08$ $77.48\pm 1.97$ $77.48\pm 1.18$ $67.45\pm 0.93$ $77.48\pm 1.67$ $78.00\pm 2.10$ $70.19\pm 4.18$ $77.87\pm 1.63$ $\textbf{78.27}\pm 3.66$
blood transfusion $79.61\pm 1.61$ $79.48\pm 1.09$ $79.21\pm 1.91$ $76.61\pm 0.88$ $79.14\pm 1.39$ $78.34\pm 1.52$ $76.27\pm 0.27$ $76.27\pm 0.27$ $\textbf{80.42}\pm 3.14$
seismic bumps $93.32\pm 0.09$ $93.30\pm 0.18$ $92.86\pm 0.34$ $93.40\pm 0.16$ $93.11\pm 0.31$ $\textbf{93.42}\pm 0.06$ $90.71\pm 5.19$ $\textbf{93.42}\pm 0.01$ $\textbf{93.42}\pm 0.01$
haberman $76.96\pm 3.31$ $76.97\pm 2.56$ $75.49\pm 2.48$ $74.02\pm 0.77$ $75.49\pm 1.49$ $76.48\pm 4.05$ $73.69\pm 0.32$ $73.53\pm 0.48$ $\textbf{77.13}\pm 2.65$
wholesale $91.02\pm 2.87$ $91.93\pm 2.19$ $90.80\pm 2.28$ $69.20\pm 1.36$ $90.80\pm 2.12$ $91.70\pm 2.18$ $69.09\pm 1.26$ $91.59\pm 1.96$ $\textbf{92.05}\pm 2.25$
liver disorder $72.76\pm 5.61$ $69.61\pm 3.65$ $72.76\pm 5.45$ $59.87\pm 2.68$ $71.50\pm 5.32$ $70.39\pm 5.06$ $66.94\pm 4.71$ $69.45\pm 4.10$ $\textbf{73.69}\pm 4.69$
wine $78.85\pm 3.82$ $\textbf{100.00}\pm 0.00$ $76.54\pm 7.81$ $69.23\pm 6.81$ $95.38\pm 3.68$ $\textbf{100.00}\pm 0.00$ $70.38\pm 6.73$ $99.62\pm 0.77$ $\textbf{100.00}\pm 0.00$
balance $99.74\pm 0.52$ $95.66\pm 1.44$ $99.39\pm 0.49$ $65.01\pm 3.22$ $99.13\pm 0.93$ $95.75\pm 0.89$ $61.89\pm 2.51$ $95.48\pm 1.49$ $\textbf{99.78}\pm 0.22$
taxpayer $\textbf{55.40}\pm 3.56$ $53.90\pm 3.25$ $53.67\pm 1.67$ $50.37\pm 0.24$ $53.97\pm 3.23$ $53.82\pm 2.85$ $51.28\pm 2.44$ $51.28\pm 2.57$ $54.65\pm 3.28$
car evaluation $95.14\pm 0.73$ $86.54\pm 1.15$ $94.42\pm 0.93$ $80.40\pm 1.18$ $92.10\pm 0.94$ $86.54\pm 1.98$ $81.15\pm 1.54$ $86.20\pm 1.56$ $\textbf{95.26}\pm 0.92$
CTG $98.38\pm 0.55$ $96.05\pm 0.79$ $98.02\pm 0.62$ $79.42\pm 1.70$ $97.34\pm 0.63$ $96.73\pm 0.98$ $77.85\pm 0.07$ $97.06\pm 0.93$ $\textbf{98.42}\pm 0.63$
US heart $84.99\pm 0.80$ $\textbf{85.35}\pm 0.46$ $84.83\pm 0.31$ $84.94\pm 0.15$ $84.94\pm 0.21$ $85.27\pm 0.41$ $84.75\pm 0.18$ $84.77\pm 0.06$ $85.21\pm 0.48$
Alzheimer rowdel $90.40\pm 1.19$ $90.12\pm 1.72$ $89.55\pm 0.67$ $86.30\pm 2.59$ $88.71\pm 3.46$ $90.40\pm 1.68$ $75.40\pm 9.18$ $89.55\pm 0.67$ $\textbf{90.68}\pm 1.38$
Alzheimer clinical $77.00\pm 2.90$ $76.67\pm 2.28$ $75.92\pm 2.53$ $70.88\pm 1.23$ $76.54\pm 3.04$ $75.88\pm 2.32$ $70.42\pm 0.49$ $76.71\pm 2.29$ $\textbf{77.42}\pm 2.17$
thyroid $92.09\pm 5.22$ $89.77\pm 5.22$ $91.63\pm 4.56$ $75.81\pm 4.05$ $92.09\pm 4.31$ $91.16\pm 3.42$ $90.23\pm 5.58$ $88.84\pm 4.00$ $\textbf{95.35}\pm 2.55$
ecoli 1 $97.33\pm 1.73$ $97.33\pm 1.73$ $97.33\pm 1.73$ $67.55\pm 2.94$ $96.44\pm 2.41$ $97.03\pm 1.88$ $68.74\pm 1.09$ $96.73\pm 2.19$ $\textbf{97.92}\pm 1.52$
ecoli 123 $91.38\pm 1.92$ $90.20\pm 3.99$ $90.78\pm 1.69$ $80.95\pm 0.58$ $90.18\pm 2.57$ $89.60\pm 3.35$ $80.95\pm 0.58$ $89.29\pm 3.29$ $\textbf{91.97}\pm 2.20$
ecoli 4 $93.15\pm 1.22$ $93.15\pm 0.75$ $91.96\pm 1.81$ $81.27\pm 4.67$ $92.55\pm 1.65$ $92.84\pm 2.58$ $90.76\pm 2.92$ $92.85\pm 1.48$ $\textbf{94.04}\pm 1.35$
ecoli 1234 $97.91\pm 1.52$ $97.62\pm 1.52$ $97.91\pm 0.73$ $91.37\pm 0.59$ $96.73\pm 1.46$ $97.03\pm 1.62$ $91.37\pm 0.59$ $97.62\pm 1.20$ $\textbf{98.21}\pm 1.46$
ecoli 5 $\textbf{99.41}\pm 0.73$ $98.82\pm 2.35$ $98.81\pm 1.10$ $78.30\pm 5.17$ $98.51\pm 0.01$ $99.11\pm 1.18$ $94.05\pm 2.50$ $98.81\pm 1.10$ $\textbf{99.41}\pm 1.18$
ecoli 678 $99.11\pm 1.19$ $99.11\pm 0.96$ $98.51\pm 1.13$ $97.32\pm 0.59$ $98.81\pm 1.28$ $\textbf{99.41}\pm 0.96$ $97.32\pm 0.59$ $99.26\pm 0.96$ $\textbf{99.41}\pm 0.96$
Average rank of accuracy 2.71 3.95 4.76 8.00 5.19 4.05 8.19 5.48 1.14
@@
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM Im-LS-$\mathfrak{U}$-QTSVM
pima 0.141 0.013 0.128 0.181 0.537 0.671 0.798 0.035 0.199
blood transfusion 0.053 0.013 0.047 0.184 0.664 0.993 1.025 0.027 0.106
seismic bumps 1.690 0.079 1.555 4.766 41.329 33.805 54.145 0.126 2.847
haberman 0.019 0.008 0.016 0.043 0.097 0.152 0.178 0.009 0.032
wholesale 0.067 0.007 0.062 0.066 0.205 0.198 0.257 0.010 0.088
liver disorder 0.038 0.007 0.034 0.048 0.099 0.121 0.057 0.011 0.047
wine 0.065 0.004 0.065 0.008 0.082 0.021 0.019 0.004 0.070
balance 0.040 0.010 0.034 0.104 0.277 0.408 0.324 0.007 0.031
taxpayer 0.150 0.012 0.140 0.123 0.417 0.480 0.475 0.043 0.156
car evaluation 0.195 0.039 0.164 1.031 3.272 3.861 5.591 0.088 0.345
CTG 3.303 0.059 3.126 2.399 9.994 11.610 2.344 0.070 4.118
US heart 2.719 0.137 2.211 10.382 59.138 72.856 103.524 0.400 3.661
Alzheimer rowdel 0.081 0.007 0.077 0.056 0.230 0.252 0.243 0.011 0.141
Alzheimer clinical 0.114 0.024 0.096 0.590 1.965 2.174 2.189 0.070 0.178
thyroid 0.022 0.005 0.020 0.020 0.072 0.071 0.054 0.006 0.042
ecoli 1 0.052 0.007 0.050 0.045 0.140 0.130 0.119 0.005 0.061
ecoli 123 0.050 0.007 0.046 0.061 0.158 0.127 0.040 0.006 0.087
ecoli 4 0.051 0.007 0.053 0.035 0.176 0.274 0.219 0.005 0.099
ecoli 1234 0.052 0.006 0.046 0.025 0.186 0.155 0.062 0.005 0.096
ecoli 5 0.053 0.007 0.048 0.031 0.172 0.198 0.246 0.005 0.094
ecoli 678 0.051 0.007 0.049 0.049 0.239 0.205 0.054 0.004 0.097
@@
For each tested model, the mean and the standard deviation of the classification accuracy on each data set are recorded in Table 3. The average rank of accuracy scores on all benchmark datasets is also calculated for each tested model. Moreover, the highest accuracy score on each benchmark dataset is highlighted. In addition, we have recorded the CPU time consumed by each tested model on all the benchmark datasets in Table 4.
@@
The observations from the results in Tables 3 and 4 can be summarized as follows.
  • Compared with all the other tested models, the proposed Im-LS-$\mathfrak{U}$-QTSVM achieves the highest mean accuracy scores on most of the benchmark datasets. Besides, it provides the highest average rank of accuracy score. In other words, the proposed model provides a better general classification accuracy than other tested models does in this experiment.
  • In general, the CPU time consumed by the Im-LS-$\mathfrak{U}$-QTSVM model is acceptable. As a least squares model, the proposed Im-LS-$\mathfrak{U}$-QTSVM does not lose efficiency when the size data set increases. Even for the relative large-scale datasets seismic bumps, US heart and CTG, the CPU time consumed by the proposed model is acceptable, and is less than that consumed by the other twin SVM models without least squares.

### v2

@@ 5.2 Results on Public Benchmark Data
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM CSSVM URKWELM FHTPSVM Im-LS-$\mathfrak{U}$-QTSVM
pima $7.74\pm 2.08 77$$48\pm 1.97 77.$$7.48\pm 1.18$ $67.45\pm 0.93$ $77.48\pm 1.67$ $78.00\pm 2.10$ $0.19\pm 4.18 77$$7.87\pm 1.63$ $76.04\pm 3.58$ $2.59\pm 2.86 77$$7.15\pm 2.01$ $\textbf{78.27}\pm 3.66$
blood transfusion $79.61\pm 1.61$ $79.48\pm 1.09$ $79.21\pm 1.91$ $76.61\pm 0.88$ $79.14\pm 1.39$ $78.34\pm 1.52$ $76.27\pm 0.27$ $76.27\pm 0.27$ $5.27\pm 6.33 77$$7.00\pm 1.17$ $78.47\pm 1.73$ $\textbf{80.42}\pm 3.14$
seismic bumps $93.32\pm 0.09$ $93.30\pm 0.18$ $92.86\pm 0.34$ $93.40\pm 0.16$ $93.11\pm 0.31$ $\textbf{93.42}\pm 0.06$ $90.71\pm 5.19$ $\textbf{93.42}\pm 0.01$ $80.03\pm 1.67$ $93.44\pm 0.06$ $93.42\pm 0.08$ $\textbf{93.42}\pm 0.01$
[…]
wholesale $91.02\pm 2.87$ $91.93\pm 2.19$ $90.80\pm 2.28$ $69.20\pm 1.36$ $90.80\pm 2.12$ $91.70\pm 2.18$ $69.09\pm 1.26$ $91.59\pm 1.96$ $91.36\pm 2.65$ $80.80\pm 4.37$ $90.80\pm 1.47$ $\textbf{92.05}\pm 2.25$
liver disorder $72.76\pm 5.61$ $69.61\pm 3.65$ $72.76\pm 5.45$ $59.87\pm 2.68$ $71.50\pm 5.32$ $70.39\pm 5.06$ $66.94\pm 4.71$ $69.45\pm 4.10$ $67.26\pm 5.84$ $58.15\pm 0.52$ $68.35\pm 4.48$ $\textbf{73.69}\pm 4.69$
wine $78.85\pm 3.82$ $\textbf{100.00}\pm 0.00$ $76.54\pm 7.81$ $69.23\pm 6.81$ $95.38\pm 3.68$ $\textbf{100.00}\pm 0.00$ $70.38\pm 6.73$ $9.62\pm 0.77 99$$9.23\pm 1.54$ $3.85\pm 4.62 99$$9.62\pm 1.09$ $\textbf{100.00}\pm 0.00$
[…]
taxpayer $\textbf{55.40}\pm 3.56$ $53.90\pm 3.25$ $53.67\pm 1.67$ $50.37\pm 0.24$ $53.97\pm 3.23$ $53.82\pm 2.85$ $51.28\pm 2.44$ $51.28\pm 2.57$ $51.58\pm 3.14$ $52.63\pm 5.12$ $52.40\pm 3.04$ $54.65\pm 3.28$
car evaluation $95.14\pm 0.73$ $86.54\pm 1.15$ $94.42\pm 0.93$ $80.40\pm 1.18$ $92.10\pm 0.94$ $86.54\pm 1.98$ $81.15\pm 1.54$ $86.20\pm 1.56$ $81.71\pm 1.84$ $73.62\pm 4.65$ $86.29\pm 1.70$ $\textbf{95.26}\pm 0.92$
CTG $98.38\pm 0.55$ $96.05\pm 0.79$ $98.02\pm 0.62$ $79.42\pm 1.70$ $97.34\pm 0.63$ $96.73\pm 0.98$ $77.85\pm 0.07$ $97.06\pm 0.93$ $96.07\pm 0.57$ $93.41\pm 1.06$ $96.14\pm 0.66$ $\textbf{98.42}\pm 0.63$
US heart $84.99\pm 0.80$ $\textbf{85.35}\pm 0.46$ $84.83\pm 0.31$ $84.94\pm 0.15$ $84.94\pm 0.21$ $85.27\pm 0.41$ $84.75\pm 0.18$ $84.77\pm 0.06$ $68.85\pm 2.07$ $82.41\pm 1.02$ $84.90\pm 0.24$ $85.21\pm 0.48$
Alzheimer rowdel $90.40\pm 1.19$ $90.12\pm 1.72$ $89.55\pm 0.67$ $6.30\pm 2.59 88$$8.71\pm 3.46$ $90.40\pm 1.68$ $75.40\pm 9.18$ $89.55\pm 0.67$ $71.03\pm 5.49$ $89.55\pm 0.67$ $89.84\pm 1.17$ $\textbf{90.68}\pm 1.38$
[…]
thyroid $92.09\pm 5.22$ $89.77\pm 5.22$ $91.63\pm 4.56$ $75.81\pm 4.05$ $92.09\pm 4.31$ $91.16\pm 3.42$ $90.23\pm 5.58$ $88.84\pm 4.00$ $90.93\pm 4.67$ $86.28\pm 4.77$ $89.07\pm 4.11$ $\textbf{95.35}\pm 2.55$
ecoli 1 $97.33\pm 1.73$ $97.33\pm 1.73$ $97.33\pm 1.73$ $67.55\pm 2.94$ $96.44\pm 2.41$ $97.03\pm 1.88$ $68.74\pm 1.09$ $96.73\pm 2.19$ $97.18\pm 2.25$ $97.17\pm 2.62$ $97.32\pm 2.06$ $\textbf{97.92}\pm 1.52$
ecoli 123 $91.38\pm 1.92$ $90.20\pm 3.99$ $90.78\pm 1.69$ $80.95\pm 0.58$ $90.18\pm 2.57$ $89.60\pm 3.35$ $80.95\pm 0.58$ $89.29\pm 3.29$ $87.79\pm 4.70$ $81.10\pm 1.48$ $89.29\pm 2.76$ $\textbf{91.97}\pm 2.20$
ecoli 4 $93.15\pm 1.22$ $93.15\pm 0.75$ $91.96\pm 1.81$ $81.27\pm 4.67$ $92.55\pm 1.65$ $92.84\pm 2.58$ $90.76\pm 2.92$ $92.85\pm 1.48$ $88.98\pm 2.71$ $85.56\pm 3.29$ $92.72\pm 2.22$ $\textbf{94.04}\pm 1.35$
ecoli 1234 $97.91\pm 1.52$ $97.62\pm 1.52$ $97.91\pm 0.73$ $91.37\pm 0.59$ $96.73\pm 1.46$ $97.03\pm 1.62$ $91.37\pm 0.59$ $97.62\pm 1.20$ $95.99\pm 2.38$ $92.11\pm 0.59$ $97.62\pm 1.64$ $\textbf{98.21}\pm 1.46$
[…]
Average rank of accuracy 3.19 4.38 5.48 10.57 6.00 4.33 10.62 6.67 8.76 9.19 6.52 1.29
@@
Dataset G-mean (mean $\pm$ standard deviation) (%)
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM CSSVM URKWELM FHTPSVM Im-LS-$\mathfrak{U}$-QTSVM
pima $74.02\pm 2.82$ $75.27\pm 3.49$ $54.78\pm 12.50$ $46.47\pm 3.38$ $72.43\pm 3.08$ $71.53\pm 4.81$ $72.39\pm 1.68$ $70.41\pm 3.91$ $74.56\pm 3.95$ $74.01\pm 0.50$ $74.01\pm 2.06$ $\textbf{75.62}\pm 2.31$
blood transfusion $66.41\pm 2.38$ $69.76\pm 2.20$ $54.14\pm 6.92$ $32.58\pm 5.55$ $64.61\pm 3.53$ $60.52\pm 3.33$ $67.87\pm 1.85$ $30.21\pm 9.56$ $67.69\pm 3.07$ $67.18\pm 3.27$ $67.18\pm 3.94$ $\textbf{79.68}\pm 1.24$
seismic bumps $55.00\pm 9.52$ $62.26\pm 4.99$ $53.73\pm 3.35$ $38.36\pm 3.29$ $49.73\pm 5.36$ $46.15\pm 4.33$ $65.63\pm 4.57$ $11.62\pm 16.13$ $\textbf{71.15}\pm 6.16$ $\textbf{71.15}\pm 1.97$ $\textbf{71.15}\pm 0.00$ $70.51\pm 6.77$
haberman $3.84\pm 3.90 66$$6.19\pm 4.36$ $44.00\pm 10.27$ $9.80\pm 12.00$ $59.86\pm 4.45$ $43.08\pm 10.04$ $64.22\pm 6.50$ $34.98\pm 6.43$ $57.26\pm 6.71$ $57.08\pm 17.06$ $57.08\pm 15.71$ $\textbf{66.41}\pm 3.39$
wholesale $89.23\pm 3.22$ $90.78\pm 3.85$ $52.39\pm 26.98$ $43.38\pm 8.60$ $87.98\pm 2.36$ $90.70\pm 3.86$ $56.46\pm 13.74$ $85.17\pm 5.20$ $91.22\pm 2.97$ $91.22\pm 4.96$ $91.22\pm 3.78$ $\textbf{91.54}\pm 3.73$
liver disorder $71.04\pm 8.11$ $65.02\pm 4.47$ $52.14\pm 6.39$ $29.22\pm 17.41$ $70.62\pm 6.95$ $65.72\pm 5.81$ $51.66\pm 8.09$ $6.06\pm 4.48 66$$96\pm 3.94 66.$$81\pm 0.00 66.$$6.81\pm 7.04$ $\textbf{72.90}\pm 5.15$
wine $90.15\pm 4.80$ $\textbf{100.00}\pm 0.00$ $32.57\pm 6.24$ $61.90\pm 10.11$ $96.31\pm 2.20$ $\textbf{100.00}\pm 0.00$ $62.79\pm 17.86$ $8.59\pm 1.72 99$$9.27\pm 1.46$ $97.87\pm 0.00$ $97.87\pm 3.14$ $\textbf{100.00}\pm 0.00$
balance $99.83\pm 0.35$ $94.08\pm 1.85$ $78.31\pm 2.78$ $54.13\pm 7.19$ $99.13\pm 0.78$ $95.58\pm 0.79$ $48.35\pm 6.69$ $94.95\pm 2.01$ $95.46\pm 2.09$ $94.95\pm 0.30$ $94.95\pm 2.19$ $\textbf{100.00}\pm 0.00$
taxpayer $53.52\pm 2.10$ $50.72\pm 1.49$ $46.23\pm 4.16$ $41.81\pm 6.53$ $52.96\pm 2.25$ $50.62\pm 2.48$ $47.40\pm 8.37$ $49.69\pm 3.10$ $50.91\pm 2.89$ $49\pm 12.06 49.49$$49\pm 3.74$ $\textbf{55.52}\pm 3.24$
car evaluation $95.14\pm 1.23$ $75.37\pm 4.04$ $58.79\pm 4.44$ $52.95\pm 2.45$ $94.38\pm 1.24$ $79.41\pm 3.56$ $82.45\pm 1.98$ $77.96\pm 3.98$ $82.78\pm 2.34$ $82.78\pm 0.63$ $82.78\pm 10.00$ $\textbf{95.35}\pm 0.74$
CTG $\textbf{97.65}\pm 1.08$ $90.59\pm 1.92$ $86.06\pm 4.91$ $61.72\pm 3.05$ $96.92\pm 0.81$ $93.44\pm 1.81$ $49.02\pm 17.29$ $95.13\pm 1.83$ $95.67\pm 0.94$ $95.49\pm 0.67$ $95.49\pm 1.58$ $97.44\pm 1.45$
US heart $52.21\pm 2.58$ $44.54\pm 3.64$ $34.38\pm 13.70$ $36.48\pm 3.34$ $50.23\pm 4.28$ $48.21\pm 6.03$ $60.49\pm 2.04$ $56.41\pm 6.30$ $6.95\pm 3.18 66$$79\pm 3.40 66.$$6.79\pm 2.53$ ${67.43}\pm 3.31$
[…]
Alzheimer clinical $77.43\pm 2.61$ $\textbf{78.43}\pm 3.43$ $56.89\pm 11.68$ $35.72\pm 4.18$ $77.52\pm 1.34$ $61.05\pm 17.09$ $53.87\pm 3.15$ $72.64\pm 4.60$ $76.85\pm 3.78$ $76.85\pm 3.49$ $76.85\pm 0.00$ $78.37\pm 2.47$
thyroid $90.23\pm 3.61$ $85.78\pm 4.60$ $82.40\pm 6.36$ $39.19\pm 21.55$ $86.07\pm 8.64$ $86.89\pm 5.95$ $82.41\pm 12.77$ $80.38\pm 7.52$ $87.38\pm 6.50$ $87.38\pm 15.70$ $87.38\pm 6.43$ $\textbf{91.41}\pm 4.26$
ecoli 1 $97.39\pm 1.19$ $97.62\pm 2.02$ $97.33\pm 2.18$ $54.17\pm 5.58$ $96.06\pm 2.35$ $97.02\pm 1.21$ $51.47\pm 2.40$ $96.64\pm 2.13$ $97.36\pm 2.37$ $97.36\pm 1.80$ $97.36\pm 4.46$ $\textbf{97.95}\pm 1.80$
ecoli 123 $88.74\pm 3.53$ $\textbf{90.24}\pm 3.45$ $43.32\pm 13.54$ $57.77\pm 8.98$ $82.25\pm 6.08$ $86.06\pm 6.16$ $47.82\pm 7.74$ $79.41\pm 7.71$ $9.88\pm 3.93 88$$41\pm 0.00 88.$$8.41\pm 11.42$ $\textbf{89.73}\pm 4.36$
ecoli 4 $75.41\pm 14.30$ $88.02\pm 7.56$ $84.73\pm 6.13$ $59.30\pm 6.12$ $78.80\pm 7.05$ $78.49\pm 4.61$ $90.09\pm 5.35$ $71.67\pm 6.63$ $87.66\pm 7.30$ $86.88\pm 2.38$ $86.88\pm 28.25$ $\textbf{91.37}\pm 4.21$
ecoli 1234 $89.83\pm 9.61$ $83.62\pm 10.94$ $57.07\pm 14.60$ $53.66\pm 10.18$ $87.22\pm 10.43$ $89.69\pm 9.73$ $60.94\pm 21.23$ $87.13\pm 13.96$ $92.40\pm 9.47$ $87.04\pm 0.00$ $87.04\pm 22.52$ $\textbf{95.91}\pm 4.68$
ecoli 5 $91.69\pm 6.79$ $89.68\pm 5.60$ $64.19\pm 19.36$ $62.50\pm 8.89$ $93.90\pm 6.87$ $91.46\pm 11.60$ $49.83\pm 21.56$ $94.18\pm 6.50$ $97.75\pm 0.96$ $93.13\pm 4.21$ $93.13\pm 30.28$ $\textbf{98.40}\pm 0.88$
ecoli 678 $8.13\pm 14.23 88$$8.28\pm 14.35$ $72.25\pm 15.76$ $38.28\pm 46.89$ $74.03\pm 38.78$ $88.28\pm 14.35$ $62.45\pm 8.63$ $88.28\pm 14.35$ $87.87\pm 14.24$ $84.95\pm 0.00$ $84.95\pm 38.67$ $\textbf{92.80}\pm 11.43$
Average rank of G-mean 4.48 5.33 10.14 11.43 6.10 7.14 8.67 8.29 3.52 4.81 4.81 1.33
Table 4: Public benchmark data G-mean results.
@@
LS-$\mathfrak{U}$-QTSVM LS-$\mathfrak{U}$-TSVM LS-QTSVM LS-TSVM $\mathfrak{U}$-QTSVM $\mathfrak{U}$-TSVM TSVM SVM CSSVM URKWELM FHTPSVM Im-LS-$\mathfrak{U}$-QTSVM
pima 0.141 0.013 0.128 0.181 0.537 0.671 0.798 0.035 0.045 1.205 0.233 0.199
blood transfusion 0.053 0.013 0.047 0.184 0.664 0.993 1.025 0.027 0.044 0.668 0.635 0.106
seismic bumps 1.690 0.079 1.555 4.766 41.329 33.805 54.145 0.126 0.484 4.742 1.940 2.847
haberman 0.019 0.008 0.016 0.043 0.097 0.152 0.178 0.009 0.013 0.112 0.071 0.032
wholesale 0.067 0.007 0.062 0.066 0.205 0.198 0.257 0.010 0.016 0.243 0.303 0.088
liver disorder 0.038 0.007 0.034 0.048 0.099 0.121 0.057 0.011 0.014 0.060 0.094 0.047
wine 0.065 0.004 0.065 0.008 0.082 0.021 0.019 0.004 0.007 0.070 0.013 0.070
balance 0.040 0.010 0.034 0.104 0.277 0.408 0.324 0.007 0.014 0.292 0.175 0.031
taxpayer 0.150 0.012 0.140 0.123 0.417 0.480 0.475 0.043 0.044 0.214 0.078 0.156
car evaluation 0.195 0.039 0.164 1.031 3.272 3.861 5.591 0.088 0.106 1.326 0.833 0.345
CTG 3.303 0.059 3.126 2.399 9.994 11.610 2.344 0.070 0.106 2.705 1.113 4.118
US heart 2.719 0.137 2.211 10.382 59.138 72.856 103.524 0.400 1.091 11.199 3.571 3.661
Alzheimer rowdel 0.081 0.007 0.077 0.056 0.230 0.252 0.243 0.011 0.014 0.091 0.231 0.141
Alzheimer clinical 0.114 0.024 0.096 0.590 1.965 2.174 2.189 0.070 0.081 1.013 0.691 0.178
thyroid 0.022 0.005 0.020 0.020 0.072 0.071 0.054 0.006 0.011 0.201 0.048 0.042
ecoli 1 0.052 0.007 0.050 0.045 0.140 0.130 0.119 0.005 0.009 0.103 0.094 0.061
ecoli 123 0.050 0.007 0.046 0.061 0.158 0.127 0.040 0.006 0.012 0.082 0.052 0.087
ecoli 4 0.051 0.007 0.053 0.035 0.176 0.274 0.219 0.005 0.010 0.142 0.023 0.099
ecoli 1234 0.052 0.006 0.046 0.025 0.186 0.155 0.062 0.005 0.010 0.093 0.252 0.096
ecoli 5 0.053 0.007 0.048 0.031 0.172 0.198 0.246 0.005 0.009 0.125 0.375 0.094
ecoli 678 0.051 0.007 0.049 0.049 0.239 0.205 0.054 0.004 0.011 0.229 0.031 0.097
@@
For each tested model, we record the mean and the standard deviation of the classification accuracy and G-mean on each data set in Tables 3 and 4, respectively. The average ranks of accuracy and G-mean scores on all benchmark datasets are also calculated for each tested model. Moreover, the highest accuracy and G-mean scores on each benchmark dataset are highlighted. In addition, we have recorded the CPU time consumed by each tested model on all the benchmark datasets in Table 5.
@@
The observations from the results in Tables 3 and 5 can be summarized as follows.
  • Compared with all the other tested models, the proposed Im-LS-$\mathfrak{U}$-QTSVM achieves the highest mean accuracy and G-mean scores on most of the benchmark datasets. Besides, it provides the highest average rank of both accuracy and G-mean score. In other words, the proposed model provides a better general classification accuracy than other tested models do in this experiment.
  • In general, the CPU time consumed by the Im-LS-$\mathfrak{U}$-QTSVM model is acceptable. As a least squares model, the proposed Im-LS-$\mathfrak{U}$-QTSVM does not lose efficiency when the size of data set increases. Even for the relatively large-scale datasets seismic bumps, US heart, and CTG, the CPU time consumed by the proposed model is acceptable, and is less than that consumed by the other twin SVM models without least squares.